Madeline's car used 15 gallons to travel 390 miles. How many gallons of gas would she need to travel 156 miles?

Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16946)

15 gallons to travel 390 miles
for 1 gallon how many miles?
390/15 = 26 miles / gallon
Car gives 26 miles /gallon
For 156 miles
156 /26 = 12/2 =6 gallons

Answer by josgarithmetic(39613)